Mathematica and You

December 1, 1994

Why Mathematica?

Mathematica is a high-level…

• interactive calculator

• 1D and 2D plotting package

• symbolic calculator

• programming language

⇒ a very effective tool

Goal of tutorial is to:

• introduce syntax

• basic concepts and programming

• ideas on how to best use it